# Reliability, and Global Privacy

### v3

@@ X Conclusion
$\hskip-1.42262pt\mathrm{P_{r}}(\bm{W}_{k}^{t,-j}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|j\notin\mathcal{U}_{k}^{t})\leq e^{\frac{c_{1}}{\sqrt{\bar{\nu}_{k}^{-j}-r_{1}}}}\mathrm{P_{r}}(\bm{W}_{k}^{t}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|j\notin\mathcal{U}_{k}^{t}),$ (74)
@@
$\mathrm{P_{r}}(\bm{W}_{k}^{t,-j}\hskip-2.84526pt\in\hskip-1.42262pt\mathcal{S})\hskip-1.42262pt\leq\hskip-1.42262pte^{\frac{c_{1}}{\sqrt{\bar{\nu}_{k}^{-j}-r_{1}}}}\mathrm{P_{r}}(\bm{W}_{k}^{t}\hskip-2.84526pt\in\hskip-1.42262pt\mathcal{S})\hskip-1.42262pt+\hskip-1.42262pt(1\hskip-1.42262pt-\hskip-1.42262ptp_{k,j})(\delta^{\prime}\hskip-2.84526pt+\hskip-1.42262pt\delta_{j}^{(2)}).$ (75)
@@
Appendix C Proof of Theorem VI.3
@@
$\mathrm{P_{r}}(\bm{W}_{k}^{t}$ $\in\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u})$
$=$ $\mathrm{P_{r}}(j\in\mathcal{U}_{k}^{t})\mathrm{P_{r}}(\bm{W}_{k}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u},j\in\mathcal{U}_{k}^{t})$
$+\mathrm{P_{r}}(j\notin\mathcal{U}_{k}^{t})\mathrm{P_{r}}(\bm{W}_{k}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u}.j\notin\mathcal{U}_{k}^{t}).$ (76)
@@
$\mathrm{P_{r}}($ $\bm{W}_{k}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u},j\in\mathcal{U}_{k}^{t})$
$\leq\delta^{\prime}+\mathrm{P_{r}}(\mathcal{E})\mathrm{P_{r}}(\bm{W}_{k}^{t}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u},j\in\mathcal{U}_{k}^{t},\mathcal{E}),$ (77)
@@
$\mathrm{P_{r}}(\bm{W}_{k}^{t}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u},i\in\mathcal{A})$
$\leq e^{\epsilon_{k,j}^{\mathcal{A},3}}\mathrm{P_{r}}(\bm{W}_{k}^{t}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u}^{\prime},i\in\mathcal{A})+\hskip-1.42262pt\delta_{j}^{(3)}.$ (78)
@@
$\hskip 14.22636pt\sup_{\hskip-14.22636pt\begin{subarray}{c}\mathcal{A}:\mathcal{A}\in\mathcal{U}_{k},j,k\in\mathcal{A}\\ \lvert\nu_{k}^{t,-j}-\bar{\nu}_{k}^{-j}\rvert\leq r_{1}\end{subarray}}\hskip 0.0pt\left\lVert\bm{W}_{k}^{t,-j}+\bm{u}+\bm{N}_{j}^{t}-\bm{W}_{k}^{t,-j}-\bm{u}^{\prime}-\bm{N}_{j}^{t}\right\rVert$
@@
$\hskip-1.42262pt\mathrm{P_{r}}(\bm{W}_{k}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u},j\in\mathcal{U}_{k}^{t})$
$\leq\delta^{\prime}+e^{\frac{c_{2}}{\sqrt{\bar{\nu}_{k}-r_{1}}}}\mathrm{P_{r}}(\bm{W}_{k}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u}^{\prime},j\in\mathcal{U}_{k}^{t})\hskip-1.42262pt+\hskip-1.42262pt\delta_{j}^{(3)}.$ (80)
@@
$\mathrm{P_{r}}(\bm{W}_{k}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u})\leq$ $e^{\frac{c_{2}}{\sqrt{\bar{\nu}_{k}-r_{1}}}}\mathrm{P_{r}}(\bm{W}_{k}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u}^{\prime})$
@@
Appendix D Proof of Theorem VI.4
@@
Consider the events such that resulting variance $\nu_{k}^{t,-j}$ lies within the event $\mathcal{E}\triangleq\{\lvert\nu_{f^{\prime}}^{t,\{k,-j\}}-\bar{\nu}_{f^{\prime}}^{t,\{k,-j\}}\rvert\leq r_{2}\}$, where $r_{2}$ is chosen such that $\mathrm{P_{r}}(\mathcal{E}^{\mathrm{c}})<\delta^{\prime}$, $\delta^{\prime}\in(0,1)$, then
@@
For each realization of $\mathcal{U}_{k}^{t}=\mathcal{A}$ and $\mathcal{V}^{t}=\mathcal{B}$, similar to (67), the privacy loss is identical to that of the Gaussian mechanism, with variance $\nu_{f^{\prime}}^{t,\{k,-j\}}$. The $\ell_{2}$-sensitivity is bounded by
$\hskip 11.38109pt\sup_{\hskip-5.69054pt\begin{subarray}{c}\mathcal{A},\mathcal{B}:\mathcal{A}\in\mathcal{U}_{k},j,k\in\mathcal{A},\\ \mathcal{B}\in[K],k\in\mathcal{B}\\ \lvert\nu_{f^{\prime}}^{t,\{k,-j\}}-\bar{\nu}_{f^{\prime}}^{t,\{k,-j\}}\rvert\leq r_{2}\end{subarray}}\hskip 0.0pt\left\lVert c_{f^{\prime},k}g_{k,j}\Delta\bm{\Theta}_{j}^{t}-c_{f^{\prime},k}g_{k,j}\bm{N}_{j}^{t}\right\rVert$
@@
with probability $1-e^{-\frac{D}{2}(\delta_{0}-\ln{(1+\delta_{0})})}$, where Chi-Square tail bound is applied in (84). Consequently, for each realization of $\mathcal{U}_{k}^{t}=\mathcal{A}$ and $\mathcal{V}^{t}=\mathcal{B}$, $\bm{F}_{f^{\prime}}^{t,\{k,-j\}}$ is $(\epsilon_{f^{\prime},\{k,j\}}^{\mathcal{A},4},\delta_{f^{\prime},\{k,j\}}^{(4)})$-differentially private, where $\epsilon_{f^{\prime},\{k,j\}}^{\mathcal{A},4}=\hskip-2.84526pt\left[2\log\left(\frac{1.25}{\delta_{f^{\prime},\{k,j\}}^{(4)}}\right)\right]^{\frac{1}{2}}\hskip-2.84526pt\frac{\lvert c_{f^{\prime},k}g_{k,j}\rvert(R+\lambda\sqrt{D(1+\delta_{0})})}{\sqrt{\nu_{f^{\prime}}^{t,\{k,-j\}}}}\hskip-2.84526pt\triangleq\frac{c_{4}}{\sqrt{\nu_{f^{\prime}}^{t,\{k,-j\}}}}$. Thus, it holds that
@@
$\leq\hskip-14.22636pt\sum_{\hskip-5.69054pt\begin{subarray}{c}\mathcal{A},\mathcal{B}:\mathcal{A}\in\mathcal{U}_{k},j,k\in\mathcal{A},\\ \mathcal{B}\in[K],k\in\mathcal{B}\\ \lvert\nu_{f^{\prime}}^{t,\{k,-j\}}-\bar{\nu}_{f^{\prime}}^{t,\{k,-j\}}\rvert\leq r_{2}\end{subarray}}\hskip-14.22636pt\mathrm{P_{r}}(\{j\in\mathcal{A}\}\cap\{k\in\mathcal{B}\})\Bigg{(}e^{\frac{c_{1}}{\sqrt{\nu_{f^{\prime}}^{t,\{k,-j\}}}}}\cdot$
@@
$\leq e^{\frac{c_{4}}{\sqrt{\bar{\nu}_{f^{\prime}}^{t,\{k,-j\}}-r_{2}}}}\cdot\hskip-14.22636pt\sum_{\hskip-5.69054pt\begin{subarray}{c}\mathcal{A},\mathcal{B}:\mathcal{A}\in\mathcal{U}_{k},j,k\in\mathcal{A},\\ \mathcal{B}\in[K],k\in\mathcal{B}\\ \lvert\nu_{f^{\prime}}^{t,\{k,-j\}}-\bar{\nu}_{f^{\prime}}^{t,\{k,-j\}}\rvert\leq r_{2}\end{subarray}}\hskip-14.22636pt\mathrm{P_{r}}(\{j\in\mathcal{A}\}\cap\{k\in\mathcal{B}\})\cdot$
@@
$\leq e^{\frac{c_{4}}{\sqrt{\bar{\nu}_{f^{\prime}}^{t,\{k,-j\}}-r_{2}}}}\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\in\mathcal{S}|\{j\in\mathcal{U}_{k}^{t}\}\cap\{k\in\mathcal{V}^{t}\},\mathcal{E})\hskip-1.42262pt$
@@
where (85c) holds since $\nu_{f^{\prime}}^{t,\{k,-j\}}>\bar{\nu}_{f^{\prime}}^{t,\{k,-j\}}-r_{2}$, (85d) is similar to (66).
@@
$\leq$ $\mathrm{P_{r}}(\mathcal{E})\bigg{(}e^{\frac{c_{4}}{\sqrt{\bar{\nu}_{f^{\prime}}^{t,\{k,-j\}}-r_{2}}}}\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\in\mathcal{S}|\{j\in\mathcal{U}_{k}^{t}\}\cap\{k\in\mathcal{V}^{t}\},\mathcal{E})$
@@
$\leq$ $e^{\frac{c_{4}}{\sqrt{\bar{\nu}_{f^{\prime}}^{t,\{k,-j\}}-r_{2}}}}\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\in\mathcal{S}|\{j\in\mathcal{U}_{k}^{t}\}\cap\{k\in\mathcal{V}^{t}\})$
@@
$\leq e^{\frac{c_{4}}{\sqrt{\bar{\nu}_{f^{\prime}}^{t,\{k,-j\}}-r_{2}}}}\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\in\mathcal{S}|\{j\notin\mathcal{U}_{k}^{t}\}\cup\{k\notin\mathcal{V}^{t}\}).$ (87)
@@
$\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t,\{k,-j\}}\in\mathcal{S})\leq e^{\frac{c_{4}}{\sqrt{\bar{\nu}_{f^{\prime}}^{t,\{k,-j\}}-r_{2}}}}\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\in\mathcal{S})$
@@
Consider the events such that resulting variance $\nu_{f^{\prime}}^{t}$ lies within the event $\mathcal{E}\triangleq\{\lvert\nu_{f^{\prime}}^{t}-\bar{\nu}_{f^{\prime}}^{t}\rvert\leq r_{2}\}$, where $r_{2}$ is chosen such that $\mathrm{P_{r}}(\mathcal{E}^{\mathrm{c}})<\delta^{\prime}$, $\delta^{\prime}\in(0,1)$, then
@@
For each realization of $\mathcal{U}_{k}^{t}=\mathcal{A}$ and $\mathcal{V}^{t}=\mathcal{B}$, similar to (67), the privacy loss is identical to that of the Gaussian mechanism, with variance $\nu_{f^{\prime}}^{t}$. The $\ell_{2}$-sensitivity is bounded by
$\hskip 11.38109pt\sup_{\hskip-5.69054pt\begin{subarray}{c}\mathcal{A},\mathcal{B}:\mathcal{B}\in[K],\mathcal{B}\cap\mathcal{V}_{f^{\prime}}=\emptyset,\\ k\in\mathcal{B},\mathcal{A}\in\mathcal{U}_{k},j,k\in\mathcal{A},\\ \lvert\nu_{f^{\prime}}^{t}-\bar{\nu}_{f^{\prime}}^{t}\rvert\leq r_{2}\end{subarray}}\hskip 0.0pt\left\lVert\sum_{k\in\mathcal{B}}c_{f^{\prime},k}g_{k,j}(\bm{u}-\bm{u}^{\prime})\right\rVert$
@@
Consequently, for each realization of $\mathcal{U}_{k}^{t}=\mathcal{A}$ and $\mathcal{V}^{t}=\mathcal{B}$, $\bm{F}_{f^{\prime}}^{t}$ is $(\epsilon_{f^{\prime},j}^{\mathcal{A},\mathcal{B},5},\delta_{f^{\prime},j}^{(5)})$-differentially private, where $\epsilon_{f^{\prime},j}^{\mathcal{A},\mathcal{B},5}=\hskip-2.84526pt\left[2\log\left(\frac{1.25}{\delta_{f^{\prime},j}^{(5)}}\right)\right]^{\frac{1}{2}}\hskip-2.84526pt\frac{2R\sum_{k=1}^{K}\lvert c_{f^{\prime},k}g_{k,j}\rvert}{\sqrt{\nu_{f^{\prime}}^{t}}}\hskip-2.84526pt\triangleq\frac{c_{5}}{\sqrt{\nu_{f^{\prime}}^{t}}}$.
@@
$=\sum_{\hskip-5.69054pt\begin{subarray}{c}\mathcal{A},\mathcal{B}:\mathcal{B}\in[K],\mathcal{B}\cap\mathcal{V}_{f^{\prime}}=\emptyset,\\ k\in\mathcal{B},\mathcal{A}\in\mathcal{U}_{k},j,k\in\mathcal{A},\\ \lvert\nu_{f^{\prime}}^{t}-\bar{\nu}_{f^{\prime}}^{t}\rvert\leq r_{2}\end{subarray}}\mathrm{P_{r}}(\mathcal{A},\mathcal{B})\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\hskip-2.84526pt\in\hskip-2.84526pt\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}\hskip-2.84526pt=\bm{u},\mathcal{J}(\mathcal{A},\mathcal{B}),\mathcal{E})$ (92a)
$\leq\hskip 0.0pt\sum_{\hskip-28.45274pt\begin{subarray}{c}\mathcal{A},\mathcal{B}:\mathcal{B}\in[K],\mathcal{B}\cap\mathcal{V}_{f^{\prime}}=\emptyset,\\ k\in\mathcal{B},\mathcal{A}\in\mathcal{U}_{k},j,k\in\mathcal{A},\\ \lvert\nu_{f^{\prime}}^{t}-\bar{\nu}_{f^{\prime}}^{t}\rvert\leq r_{2}\end{subarray}}\mathrm{P_{r}}(\mathcal{A},\mathcal{B})\big{(}e^{\epsilon_{f^{\prime},j}^{\mathcal{A},\mathcal{B},5}}\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\hskip-2.84526pt\in\hskip-2.84526pt\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}\hskip-2.84526pt=\bm{u}^{\prime},\mathcal{J}(\mathcal{A},\mathcal{B}),\mathcal{E})$
@@
$\leq e^{\frac{c_{5}}{\sqrt{\bar{\nu}_{f^{\prime}}^{t}-r_{2}}}}\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\hskip-2.84526pt\in\hskip-2.84526pt\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}\hskip-2.84526pt=\bm{u}^{\prime},\mathcal{J},\mathcal{E})+\delta_{f^{\prime},j}^{(5)}.$ (92c)

### v4

@@ VIII Conclusion
$\hskip-1.42262pt\mathrm{P_{r}}(\bm{S}_{k}^{t,-j}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|j\notin\mathcal{U}_{k}^{t})\leq e^{\frac{c_{1}}{\sqrt{\bar{\nu}_{k}^{-j}-r_{1}}}}\mathrm{P_{r}}(\bm{S}_{k}^{t}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|j\notin\mathcal{U}_{k}^{t}),$ (62)
@@
$\mathrm{P_{r}}(\bm{S}_{k}^{t,-j}\hskip-2.84526pt\in\hskip-1.42262pt\mathcal{S})\hskip-1.42262pt\leq\hskip-1.42262pte^{\frac{c_{1}}{\sqrt{\bar{\nu}_{k}^{-j}-r_{1}}}}\mathrm{P_{r}}(\bm{S}_{k}^{t}\hskip-2.84526pt\in\hskip-1.42262pt\mathcal{S})\hskip-1.42262pt+\hskip-1.42262pt(1\hskip-1.42262pt-\hskip-1.42262ptp_{k,j})(\delta^{\prime}\hskip-2.84526pt+\hskip-1.42262pt\delta_{j}^{(2)}).$ (63)
@@
Appendix C Proof of Theorem V.3
@@
$\mathrm{P_{r}}(\bm{S}_{k}^{t}$ $\in\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u})$
$=$ $\mathrm{P_{r}}(j\in\mathcal{U}_{k}^{t})\mathrm{P_{r}}(\bm{S}_{k}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u},j\in\mathcal{U}_{k}^{t})$
$+\mathrm{P_{r}}(j\notin\mathcal{U}_{k}^{t})\mathrm{P_{r}}(\bm{S}_{k}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u}.j\notin\mathcal{U}_{k}^{t}).$ (64)
@@
$\mathrm{P_{r}}($ $\bm{S}_{k}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u},j\in\mathcal{U}_{k}^{t})$
$\leq\delta^{\prime}+\mathrm{P_{r}}(\mathcal{E})\mathrm{P_{r}}(\bm{S}_{k}^{t}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u},j\in\mathcal{U}_{k}^{t},\mathcal{E}),$ (65)
@@
$\mathrm{P_{r}}(\bm{S}_{k}^{t}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u},i\in\mathcal{A})$
$\leq e^{\epsilon_{k,j}^{\mathcal{A},3}}\mathrm{P_{r}}(\bm{S}_{k}^{t}\hskip-1.42262pt\in\hskip-2.84526pt\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u}^{\prime},i\in\mathcal{A})+\hskip-1.42262pt\delta_{j}^{(3)}.$ (66)
@@
$\hskip 14.22636pt\sup_{\hskip-14.22636pt\begin{subarray}{c}\mathcal{A}:\mathcal{A}\in\mathcal{U}_{k},j,k\in\mathcal{A}\\ \lvert\nu_{k}^{t,-j}-\bar{\nu}_{k}^{-j}\rvert\leq r_{2}\end{subarray}}\hskip 0.0pt\left\lVert\bm{S}_{k}^{t,-j}+\bm{u}+\bm{N}_{j}^{t}-\bm{S}_{k}^{t,-j}-\bm{u}^{\prime}-\bm{N}_{j}^{t}\right\rVert$
@@
$\hskip-1.42262pt\mathrm{P_{r}}(\bm{S}_{k}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u},j\in\mathcal{U}_{k}^{t})$
$\leq\delta^{\prime}+e^{\frac{c_{2}}{\sqrt{\bar{\nu}_{k}-r_{2}}}}\mathrm{P_{r}}(\bm{S}_{k}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u}^{\prime},j\in\mathcal{U}_{k}^{t})\hskip-1.42262pt+\hskip-1.42262pt\delta_{j}^{(3)}.$ (68)
@@
$\mathrm{P_{r}}(\bm{S}_{k}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u})\leq$ $e^{\frac{c_{2}}{\sqrt{\bar{\nu}_{k}-r_{2}}}}\mathrm{P_{r}}(\bm{S}_{k}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u}^{\prime})$
@@
Appendix D Proof of Theorem V.4
@@
Consider the events such that resulting variance $\nu_{k}^{t,-j}$ lies within the event $\mathcal{E}\triangleq\{\lvert\nu_{F_{f^{\prime}}}^{t,\{k,-j\}}-\bar{\nu}_{F_{f^{\prime}}}^{t,\{k,-j\}}\rvert\leq r_{3}\}$, where $r_{3}$ is chosen such that $\mathrm{P_{r}}(\mathcal{E}^{\mathrm{c}})<\delta^{\prime}$, $\delta^{\prime}\in(0,1)$, then
@@
For each realization of $\mathcal{U}_{k}^{t}=\mathcal{A}$ and $\mathcal{V}^{t}=\mathcal{B}$, similar to (55), the privacy loss is identical to that of the Gaussian mechanism, with variance $\nu_{F_{f^{\prime}}}^{t,\{k,-j\}}$. The $\ell_{2}$-sensitivity is bounded by
$\hskip 11.38109pt\sup_{\hskip-5.69054pt\begin{subarray}{c}\mathcal{A},\mathcal{B}:\mathcal{A}\in\mathcal{U}_{k},j,k\in\mathcal{A},\\ \mathcal{B}\in[K],k\in\mathcal{B}\\ \lvert\nu_{F_{f^{\prime}}}^{t,\{k,-j\}}-\bar{\nu}_{F_{f^{\prime}}}^{t,\{k,-j\}}\rvert\leq r_{3}\end{subarray}}\hskip 0.0pt\left\lVert c_{f^{\prime},k}g_{k,j}\Delta\bm{\Theta}_{j}^{t}-c_{f^{\prime},k}g_{k,j}\bm{N}_{j}^{t}\right\rVert$
@@
with probability $1-e^{-\frac{D}{2}(\delta_{0}-\ln{(1+\delta_{0})})}$, where Chi-Square tail bound is applied in (72). Consequently, for each realization of $\mathcal{U}_{k}^{t}=\mathcal{A}$ and $\mathcal{V}^{t}=\mathcal{B}$, $\bm{F}_{f^{\prime}}^{t,\{k,-j\}}$ is $(\epsilon_{f^{\prime},\{k,j\}}^{\mathcal{A},4},\delta_{f^{\prime},\{k,j\}}^{(4)})$-differentially private, where $\epsilon_{f^{\prime},\{k,j\}}^{\mathcal{A},4}=\hskip-2.84526pt\left[2\log\left(\frac{1.25}{\delta_{f^{\prime},\{k,j\}}^{(4)}}\right)\right]^{\frac{1}{2}}\hskip-2.84526pt\frac{\lvert c_{f^{\prime},k}g_{k,j}\rvert(R+\lambda\sqrt{D(1+\delta_{0})})}{\sqrt{\nu_{F_{f^{\prime}}}^{t,\{k,-j\}}}}\hskip-2.84526pt\triangleq\frac{c_{3}}{\sqrt{\nu_{F_{f^{\prime}}}^{t,\{k,-j\}}}}$. Thus, it holds that
@@
$\leq\hskip-14.22636pt\sum_{\hskip-5.69054pt\begin{subarray}{c}\mathcal{A},\mathcal{B}:\mathcal{A}\in\mathcal{U}_{k},j,k\in\mathcal{A},\\ \mathcal{B}\in[K],k\in\mathcal{B}\\ \lvert\nu_{F_{f^{\prime}}}^{t,\{k,-j\}}-\bar{\nu}_{F_{f^{\prime}}}^{t,\{k,-j\}}\rvert\leq r_{3}\end{subarray}}\hskip-14.22636pt\mathrm{P_{r}}(\{j\in\mathcal{A}\}\cap\{k\in\mathcal{B}\})\Bigg(e^{\frac{c_{1}}{\sqrt{\nu_{F_{f^{\prime}}}^{t,\{k,-j\}}}}}\cdot$
@@
$\leq e^{\frac{c_{3}}{\sqrt{\bar{\nu}_{F_{f^{\prime}}}^{t,\{k,-j\}}-r_{3}}}}\cdot\hskip-14.22636pt\sum_{\hskip-5.69054pt\begin{subarray}{c}\mathcal{A},\mathcal{B}:\mathcal{A}\in\mathcal{U}_{k},j,k\in\mathcal{A},\\ \mathcal{B}\in[K],k\in\mathcal{B}\\ \lvert\nu_{F_{f^{\prime}}}^{t,\{k,-j\}}-\bar{\nu}_{F_{f^{\prime}}}^{t,\{k,-j\}}\rvert\leq r_{3}\end{subarray}}\hskip-14.22636pt\mathrm{P_{r}}(\{j\in\mathcal{A}\}\cap\{k\in\mathcal{B}\})\cdot$
@@
$\leq e^{\frac{c_{3}}{\sqrt{\bar{\nu}_{F_{f^{\prime}}}^{t,\{k,-j\}}-r_{3}}}}\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\in\mathcal{S}|\{j\in\mathcal{U}_{k}^{t}\}\cap\{k\in\mathcal{V}^{t}\},\mathcal{E})\hskip-1.42262pt$
@@
where (73c) holds since $\nu_{F_{f^{\prime}}}^{t,\{k,-j\}}>\bar{\nu}_{F_{f^{\prime}}}^{t,\{k,-j\}}-r_{3}$, (73d) is similar to (54).
@@
$\leq$ $\mathrm{P_{r}}(\mathcal{E})\bigg(e^{\frac{c_{3}}{\sqrt{\bar{\nu}_{F_{f^{\prime}}}^{t,\{k,-j\}}-r_{3}}}}\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\in\mathcal{S}|\{j\in\mathcal{U}_{k}^{t}\}\cap\{k\in\mathcal{V}^{t}\},\mathcal{E})$
@@
$\leq$ $e^{\frac{c_{3}}{\sqrt{\bar{\nu}_{F_{f^{\prime}}}^{t,\{k,-j\}}-r_{3}}}}\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\in\mathcal{S}|\{j\in\mathcal{U}_{k}^{t}\}\cap\{k\in\mathcal{V}^{t}\})$
@@
$\leq e^{\frac{c_{3}}{\sqrt{\bar{\nu}_{F_{f^{\prime}}}^{t,\{k,-j\}}-r_{3}}}}\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\in\mathcal{S}|\{j\notin\mathcal{U}_{k}^{t}\}\cup\{k\notin\mathcal{V}^{t}\}).$ (75)
@@
$\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t,\{k,-j\}}\in\mathcal{S})\leq e^{\frac{c_{3}}{\sqrt{\bar{\nu}_{F_{f^{\prime}}}^{t,\{k,-j\}}-r_{3}}}}\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\in\mathcal{S})$
@@
Consider the events such that resulting variance $\nu_{F_{f^{\prime}}}^{t}$ lies within the event $\mathcal{E}\triangleq\{\lvert\nu_{F_{f^{\prime}}}^{t}-\bar{\nu}_{F_{f^{\prime}}}^{t}\rvert\leq r_{4}\}$, where $r_{4}$ is chosen such that $\mathrm{P_{r}}(\mathcal{E}^{\mathrm{c}})<\delta^{\prime}$, $\delta^{\prime}\in(0,1)$, then
@@
For each realization of $\mathcal{U}_{k}^{t}=\mathcal{A}$ and $\mathcal{V}^{t}=\mathcal{B}$, similar to (55), the privacy loss is identical to that of the Gaussian mechanism, with variance $\nu_{F_{f^{\prime}}}^{t}$. The $\ell_{2}$-sensitivity is bounded by
$\hskip 11.38109pt\sup_{\hskip-5.69054pt\begin{subarray}{c}\mathcal{A},\mathcal{B}:\mathcal{B}\in[K],\mathcal{B}\cap\mathcal{V}_{f^{\prime}}=\emptyset,\\ k\in\mathcal{B},\mathcal{A}\in\mathcal{U}_{k},j,k\in\mathcal{A},\\ \lvert\nu_{F_{f^{\prime}}}^{t}-\bar{\nu}_{F_{f^{\prime}}}^{t}\rvert\leq r_{4}\end{subarray}}\hskip 0.0pt\left\lVert\sum_{k\in\mathcal{B}}c_{f^{\prime},k}g_{k,j}(\bm{u}-\bm{u}^{\prime})\right\rVert$
@@
Consequently, for each realization of $\mathcal{U}_{k}^{t}=\mathcal{A}$ and $\mathcal{V}^{t}=\mathcal{B}$, $\bm{F}_{f^{\prime}}^{t}$ is $(\epsilon_{f^{\prime},j}^{\mathcal{A},\mathcal{B},5},\delta_{f^{\prime},j}^{(5)})$-differentially private, where $\epsilon_{f^{\prime},j}^{\mathcal{A},\mathcal{B},5}=\hskip-2.84526pt\left[2\log\left(\frac{1.25}{\delta_{f^{\prime},j}^{(5)}}\right)\right]^{\frac{1}{2}}\hskip-2.84526pt\frac{2R\sum_{k=1}^{K}\lvert c_{f^{\prime},k}g_{k,j}\rvert}{\sqrt{\nu_{F_{f^{\prime}}}^{t}}}\hskip-2.84526pt\triangleq\frac{c_{4}}{\sqrt{\nu_{F_{f^{\prime}}}^{t}}}$.
@@
$=\sum_{\hskip-5.69054pt\begin{subarray}{c}\mathcal{A},\mathcal{B}:\mathcal{B}\in[K],\mathcal{B}\cap\mathcal{V}_{f^{\prime}}=\emptyset,\\ k\in\mathcal{B},\mathcal{A}\in\mathcal{U}_{k},j,k\in\mathcal{A},\\ \lvert\nu_{F_{f^{\prime}}}^{t}-\bar{\nu}_{F_{f^{\prime}}}^{t}\rvert\leq r_{4}\end{subarray}}\mathrm{P_{r}}(\mathcal{A},\mathcal{B})\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\hskip-2.84526pt\in\hskip-2.84526pt\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}\hskip-2.84526pt=\bm{u},\mathcal{J}(\mathcal{A},\mathcal{B}),\mathcal{E})$ (80a)
$\leq\hskip 0.0pt\sum_{\hskip-28.45274pt\begin{subarray}{c}\mathcal{A},\mathcal{B}:\mathcal{B}\in[K],\mathcal{B}\cap\mathcal{V}_{f^{\prime}}=\emptyset,\\ k\in\mathcal{B},\mathcal{A}\in\mathcal{U}_{k},j,k\in\mathcal{A},\\ \lvert\nu_{F_{f^{\prime}}}^{t}-\bar{\nu}_{F_{f^{\prime}}}^{t}\rvert\leq r_{4}\end{subarray}}\mathrm{P_{r}}(\mathcal{A},\mathcal{B})\big(e^{\epsilon_{f^{\prime},j}^{\mathcal{A},\mathcal{B},5}}\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\hskip-2.84526pt\in\hskip-2.84526pt\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}\hskip-2.84526pt=\bm{u}^{\prime},\mathcal{J}(\mathcal{A},\mathcal{B}),\mathcal{E})$
@@
$\leq e^{\frac{c_{4}}{\sqrt{\bar{\nu}_{F_{f^{\prime}}}^{t}-r_{4}}}}\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\hskip-2.84526pt\in\hskip-2.84526pt\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}\hskip-2.84526pt=\bm{u}^{\prime},\mathcal{J},\mathcal{E})+\delta_{f^{\prime},j}^{(5)}.$ (80c)